# Vanilla Bayesian Optimization Performs Great in High Dimensions

### 25 Feb 2024 v3

@@ C.3 The Boundary Issue through the Lens of Locality
Lastly, we demonstrate the drastically different behaviors that may arise even in low-dimensional problems, such as the non-embedded variant of Hartmann (6D). Since our enhancement is applicable to low-dimensional problems as well, we demonstrate its ability to exhibit a well-calibrated exploration-exploitation trade-off. We demonstrate the performance of MLE, conventional (Balandat et al., 2020), low-lengthscale MAP ($\ell_{i}\sim\Gamma(3,6)$) and our method in Fig. 9
Figure 9: (left) Regret performance of MLE (red), a short lengthscale prior (green) and our enhancement of the lengthscale priors (blue) across 10 repetitions. Conventional MAP and our meethod perform comparably in final performance, whereas MLE is non-competitive. (right) Average distance to the incumbent for conventional MAP, MLE and our method per iteration, averaged over 50 repetitions. Conventional MAP exemplifies the locality issue, as it almost exclusively queries in close proximity to the incumbent. MLE exhibits very exploratory behavior, whereas our method displays a balance between exploration and exploitation, as visualized by the wide error bars.
While conventional MAP and our method achieve almost identical final performance, we now consider the degree to which the locality issue is present for each method by considering the minimal distance to a preceding data point for each query throughout the optimization round. Fig. 9 visualizes this, and we see that conventional MAP almost never makes an exploratory query after the DoE phase, and takes substantially smaller local steps than our enhancement. While final performance is not negatively impacted, Fig. 9 reveals that conventional MAP is effectively performing an initial random search followed by 140 iterations of local search. While potent, the high complexity of the model effectively disables the global, exploratory component of the BO algorithm.
Appendix D Ablation Studies
@@
We find that generally, Vanilla BO is highly performant across almost an order of magnitude of complexity, as displayed in Fig. 14, and very robust to changes in uncertainty of the complexity, as shown in Figs. 12 and 13. Assuming too low complexity, however, can occasionally be detrimental to performance, as displayed in Fig 17.
D.1 Conventional MAP and MLE
First, we show the performance of our enhancement to vanilla BO compared to MLE and MAP using the default BoTorch (Balandat et al., 2020) priors ($\ell_{i}\sim\Gamma(3,6)$) on the hyperparameters, which match the description given on conventional priors in Sec. 2.2. In the higher-dimensional tasks, MLE hyperparameter fitting is very unstable, due to the large number of hyperparameters in relation to the number of observations. Thus, full repetitions of MLE-estimated hyperparameters are not available for many tasks.
@@ D.1 Conventional MAP and MLE
Figure 10: Average log regret of uncertainty ablations on Levy (4D) and Hartmann (6D) synthetic test functions of varying ambient dimensionality for our method (blue), MLE-estimated hyperparameters (red) and hyperparameters fit through BoTorch default hyperpriors ($\ell_{i}\sim\Gamma(3,6)$, green) across 10 repetitions (20 for the default algorithm in blue). $\ell_{i}\sim\Gamma(3,6)$ priors perform well on the Hartmann tasks, but searches extremely locally around the best found point from DoE. MLE runs into errors related to fitting instability early on for a majority of all tasks.
Figure 11: Average minimal observed value on MOPTA, Lasso-DNA and SVM for our method (blue), MLE-estimated hyperparameters (red) and hyperparameters fit through BoTorch default hyperpriors ($\ell_{i}\sim\Gamma(3,6)$, green) across 10 repetitions (20 for the default algorithm in blue). $\ell_{i}\sim\Gamma(3,6)$ priors perform well on MOPTA, and MLE runs into errors related to fitting instability early on for all tasks.
D.2 Fixed Complexity - Varied uncertainty
@@
D.3 Varied Complexity - Fixed uncertainty
@@
Figure 14: Average log regret of Vanilla BO with higher assumed complexity (red) and lower (green) on Levy (4D) and Hartmann (6D) synthetic test functions of varying ambient dimensionality across 10 repetitions (20 for the default algorithm in blue). All instantiations perform comparably, with the low-complexity variant (green) performing slightly worse on average.
Figure 15: Average minimal observed value on MOPTA, Lasso-DNA and SVM for Vanilla BO with higher assumed complexity (red) and lower (green) across 10 repetitions (20 for the default algorithm in blue). Assuming a substantially higher complexity may have detrimental impact on performance, as demonstrated in SVM for the instantiation with the largest $\mu_{0}$.
D.4 Learning the signal variance
In our method in Sec. 5, we propose to not learn the signal variance to safeguard against possible degeneracies related to the active acquisition of data. We ablate the choice here, and demonstrate that, while the performance of a model with a learned signal variance is marginally less stable, the performance is comparable to one where the signal variance is not learned. Moreover, we visualize the signal variance that is learned for a model with conventional (Balandat et al., 2020) $\ell_{i}\sim\Gamma(3,6)$ priors, as well as the learned signal variance for our ablated variant with scaled lengthscales.
Figure 16: Average log regret of Vanilla BO with a learned signal variance (red) and the default algorithm (blue) on Levy (4D) and Hartmann (6D) synthetic test functions of varying ambient dimensionality across 10 repetitions (20 for the default algorithm in blue). The learned signal variance variant performs comparably on some tasks, but displays worse consistency across tasks.
Figure 17: Average minimal observed value of Vanilla BO with a learned signal variance (red) and the default algorithm (blue) on MOPTA, Lasso-DNA and SVM across 10 repetitions (20 for the default algorithm in blue). The learned signal variance variant performs comparably on some tasks, but displays worse consistency across tasks.

### 26 Jun 2024 v4

@@ C.3 The Boundary Issue through the Lens of Locality
Lastly, we demonstrate the drastically different behaviors that may arise even in low-dimensional problems, such as the non-embedded variant of Hartmann (6D). Since our enhancement is applicable to low-dimensional problems as well, we demonstrate its ability to exhibit a well-calibrated exploration-exploitation trade-off. We demonstrate the performance of MLE, conventional (Balandat et al., 2020), low-lengthscale MAP ($\ell_{i}\sim\Gamma(3,6)$) and the DSP in Fig. 11
Figure 11: (left) Regret performance a short, conventional lengthscale prior (yellow) and the DSP (blue) across 10 repetitions. Conventional MAP and DSP perform comparably in final performance. (right) Average distance to the incumbent for conventional MAP and DSP per iteration, averaged over 20 repetitions. Conventional MAP exemplifies the locality issue, as it almost exclusively queries in close proximity to the incumbent, whereas DSP displays a balance between exploration and exploitation, as visualized by the wide error bars.
While conventional MAP and DSP achieve almost identical final performance, we now consider the degree to which the locality issue is present for each method by considering the minimal distance to a preceding data point for each query throughout the optimization round. Fig. 11 visualizes this, and we see that conventional MAP almost never makes an exploratory query after the DoE phase, and takes substantially smaller local steps than our enhancement. While final performance is not negatively impacted, Fig. 11 reveals that conventional MAP is effectively performing an initial random search followed by 140 iterations of local search. While potent, the high complexity of the model effectively disables the global, exploratory component of the BO algorithm.
Appendix D Low-budget Performance
We display the performance of the DSP on the real-world benchmarks from Sec.6. In Fig. 12, we observe that for all real-world tasks, the DSP provides the strongest initial performance. DSP rapidly improves after initialization, which indicates potent inference ability. Notably, the DSP does not rely on identifying a subset of active variables, yet provides superior performance to the alternatives.
Figure 12: Best observed value of all baselines on five real-world tasks from of various domains across 20 repetitions (10 for SAASBO) and Humanoid. The DSP performs best across all tasks for a low-budget settings, and offers a pronounced initial jump in performance after DoE.
Appendix E Ablation Studies
@@
We find that generally, Vanilla BO is highly performant across almost an order of magnitude of complexity, as displayed in Fig. 17, and very robust to changes in uncertainty of the complexity, as shown in Figs. 15 and 16. Assuming too low complexity, however, can occasionally be detrimental to performance, as displayed in Fig 22.
E.1 Conventional MAP and MLE
First, we show the performance of our enhancement to vanilla BO compared to MLE and MAP using the default BoTorch (Balandat et al., 2020) priors ($\ell_{i}\sim\Gamma(3,6)$) on the hyperparameters, which match the description given on conventional priors in Sec. 2.2. In the higher-dimensional tasks, MLE hyperparameter fitting is very unstable, due to the large number of hyperparameters in relation to the number of observations. Thus, full repetitions of MLE-estimated hyperparameters are not available for many tasks. We run a stabilized variant of MLE fitting as a complement, where the previous iteration’s hyperparameters are in case of a crash during model fitting. This variant achieves competitive performance on many tasks, further demonstrating the capabilities of Vanilla Bayesian optimization fro high-dimensional tasks.
@@ E.1 Conventional MAP and MLE
Figure 13: Average log regret of uncertainty ablations on Levy (4D) and Hartmann (6D) synthetic test functions of varying ambient dimensionality for DSP (blue), MLE-estimated hyperparameters (red), stabilized MLE-estimated hyperparameters (purple) and hyperparameters fit through BoTorch default hyperpriors ($\ell_{i}\sim\Gamma(3,6)$, green) across 10 repetitions (20 for the default algorithm in blue). $\ell_{i}\sim\Gamma(3,6)$ priors perform well on the Hartmann tasks, but searches extremely locally around the best found point from DoE. MLE runs into errors related to fitting instability early on for a majority of all tasks.
Figure 14: Average minimal observed value on MOPTA, Lasso-DNA and SVM for DSP (blue), MLE-estimated hyperparameters (red), stabilized MLE-estimated hyperparameters (purple) and hyperparameters fit through BoTorch default hyperpriors ($\ell_{i}\sim\Gamma(3,6)$, green) across 10 repetitions (20 for the default algorithm in blue). $\ell_{i}\sim\Gamma(3,6)$ priors perform well on MOPTA, and MLE runs into errors related to fitting instability early on for all tasks.
E.2 Fixed Complexity - Varied uncertainty
@@
E.3 Varied Complexity - Fixed uncertainty
@@
Figure 17: Average log regret of DSP with higher assumed complexity (red) and lower (green) on Levy (4D) and Hartmann (6D) synthetic test functions of varying ambient dimensionality across 10 repetitions (20 for the default algorithm in blue). All instantiations perform comparably, with the low-complexity variant (green) performing slightly worse on average.
Figure 18: Average minimal observed value on MOPTA, Lasso-DNA and SVM for DSP with higher assumed complexity (red) and lower (green) across 10 repetitions (20 for the default algorithm in blue). Assuming a substantially higher complexity may have detrimental impact on performance, as demonstrated in SVM for the instantiation with the largest $\mu_{0}$.
Appendix F Learning the signal variance
In our method in Sec. 5, we propose to not learn the signal variance to safeguard against possible degeneracies related to the active acquisition of data. We ablate the choice here, and demonstrate that, while the performance of a model with a learned signal variance is marginally less stable, the performance is comparable to one where the signal variance is not learned. In Fig. 19, we display how the outputscale hyperparameter shrinks throughout optimization for models with a learned outputscale. Notably, the fixed outputscale variants (red, blue) do not display constant outputscales due to varying data normalization over time. Moreover, in Fig. 20, we show the exaggerated local search-like behavior that follows from either a learned outputscale or conventional lengthscale prior, as measured by the average distance between the current best point and the upcoming query. The initial samples of high variance and large distance between queries indicates the initial design of experiments via SOBOL sampling.
F.1 Learning $\sigma_{f}^{2}$
Figure 19: Outputscale shrinkage on MOPTA08 and SVM for BO (mean and one standard deviation) using four types of models employing either the DSP or conventional MAP, and either fixed or learned outputscale. Learned outputscale results in very pronounced shrinkage over time, which leads to increased exploitative behavior.
Figure 20: Average distance from current incumbent (mean and 0.25 standard deviation) on MOPTA08 and SVM for BO using four types of models employing either the DSP or conventional MAP, and either fixed or learned outputscale. Learned outputscale or a conventional prior leads to very local optimization, as the algorithm almost exclusively stays within very close proximity to the current best value.
We empirically demonstrate the behavior of the model when learning the signal variance for two tasks, MOPTA08 and SVM, for four types of models throughout the BO loop: DSP with fixed $\sigma_{f}^{2}$, DSP with learned $\sigma_{f}^{2}$, and a conventional $\Gamma(3,6)$ prior with and without learned $\sigma_{f}^{2}$.
F.2 Performance Ablation on Learned $\sigma_{f}^{2}$
We display the performance of the DSP with and without learned signal variance for a subset of the synthetic and real-world tasks presented in the paper. While the performance is occasionally competitve, the performance is substantially less stable than for the fixed outputscale variant.
Figure 21: Average log regret of Vanilla BO with a learned signal variance (yellow) and the default algorithm (blue) on Levy (4D) and Hartmann (6D) synthetic test functions of varying ambient dimensionality across 10 repetitions (20 for the default algorithm in blue). The learned signal variance variant performs comparably on some tasks, but displays worse consistency across tasks.
Figure 22: Average minimal observed value of Vanilla BO with a learned signal variance (yellow) and the default algorithm (blue) on MOPTA, Lasso-DNA and SVM across 10 repetitions (20 for the default algorithm in blue). The learned signal variance variant performs comparably on some tasks, but displays worse consistency across tasks.